# Veri-Car: Towards Open-world Vehicle Information Retrieval

### v3

@@ 7.2 License Plate Detection
Veri-Car uses an open source YOLOv5 model fine-tuned on the LPD task to detect and crop the license plates from car images [5]. The model can be further fine-tuned on a specific dataset. However, due to the high accuracy of the original model, our interest in maintaining generalizability, and the lack of appropriate fine-tuning datasets for multinational LPD, we have decided to use the model out of the box.
@@ 7.3 License Plate Recognition
For LPR, Veri-Car uses an open source TrOCR model. This model was first fine-tuned on the IAM handwriting dataset as it provided a good warm start relative to the pre-trained only model which was trained on a combination of textlines from PDF documents and synthetically generated handwritten text images [6, 7, 8]. Additional fine-tuning was performed on 35771 generated synthetic license plate images. We will discuss the process for generating synthetic license plate images in Section 7.5.2.
@@ 7.6.2 License Plate Recognition
The results in terms of percentage correct and CER are shown in Tables 12 and 13 respectively. As expected, the fine-tuned TrOCR model performs well on the synthetic license plates and less well on the CCPD license plates, while Tesseract fails on both. It is worth noting that in some subsets of the CCPD data, the model trained only on synthetic data performs decently, demonstrating the power of synthetic data. The TrOCR with CCPD model and the TrOCR with CCPD + Synthetic model perform very well on CCPD license plates, with the latter slightly edging out the former in most cases likely due to the larger amount of data. Given fine-tuning on just 1% of the data, however, TrOCR with CCPD is able to achieve nearly comparable performance, demonstrating that pre-training with synthetic data is a viable option in cases where there are few country specific labels. The TrOCR with CCPD + Synthetic mdel also performs well on just the synthetic data while the performance of TrOCR with CCPD in this case begins to drop off, indicating that during the fine-tuning process the model begins to learn more country specific information. Finally, the results also show that an OCR without LPR specific training, such as Tesseract, fails in the LPR setting.
@@ 7.6.3 End-to-end Evaluation
The end-to-end results shown in Tables 14 and 15 depict RPnet’s failure to capture license plates from countries on which it was not trained, even when the data has been filtered to include only license plates which fit the pattern of the training data. On the other hand, Veri-Car’s YOLO+TrOCR model, which is fine-tuned on synthetic data only, offers stronger generalizability, outperforming RPnet by a wide margin on the Stanford Cars dataset which includes license plates from a range of countries. The results further underscore the effectiveness of Veri-Car’s two additional TrOCR variants, both of which often surpass PRNet on the CCPD data. This shows that synthetic data can be valuable not only in a multinational context, but also when applied during pre-training for a country-specific detector. Similar to the LPR results, TrOCR with CCPD + Synthetic slightly outperforms TrOCR with CCPD, likely due to the increased volume of training data. Notably, within the green subset of CCPD plates, which represent images of green license plates while the training data contains only blue plates, TrOCR with CCPD significantly outperforms all other models. This suggests that the use of synthetic data is especially effective in accurately recognizing license plate numbers, despite the color variation in these plates.
@@ 8 Full System Evaluation
In this section we evaluate the effectiveness of the system using a subset of 102 samples of the Stanford Cars dataset [55]. We use the metrics already presented in each component of Veri-Car. This subset is a subset of the one we used for evaluating License Plate Recognition in the previous section. We can’t evaluate on the same set because it contains images our retrieval model was directly trained on. Our results are summarized in Table 16. This new dataset includes 86 images of cars with labels seen during training and 16 images with unseen labels, never seen during training. The results in Table 16 summarize the performance of the each component of Veri-Car for all images corresponding to seen label instances. In this section we also present the results of the OOD detection module on the unseen set of images and summarize the overall performance of the system in correctly classifying all labels of an image on both the seen and unseen sets independently.
@@
Model Metric Result Color Retrieval mAP@R 83.59 Precision@1 89.53 Make, Type, Model, Year Retrieval mAP@R 91.05 Precision@1 94.18 License Plate Recognition Percent Correct 28.43 CER 21.50 License Plate Recognition (CER <0.2) Percent Correct 49.15 CER 7.71
8.1 Make, Type, Model, Year and Color Retrieval
On one hand, our model achieved a mAP@R of 83.59 and 91.05 at color and make, type, model and year retrieval respectively in this specific subset as it can be seen in Table 16. On the other hand, Veri-Car obtains a Precision@1 of 89.53 and 94.18 on color and make, type, model and year retrieval. Based on the results, it appears that the subset is easier for color since we gain 0.75% in Precision@1. However, the subset is harder for make, type, model and year retrieval since we lost 1.74% in Precision@1.
@@ 8.3 OOD Detection
The OOD detection model results are reported in Table 17. Out of the 102 images in the test set, 16 were not seen by the model during training. The OOD detection model has an FPR95 of 37.50% and AUROC of 91.00%, highlighting increased difficulty of this specific subset of the dataset. In this case we have increase of 6.89% in FPR95 and a reduction of 1.12% in AUROC.
@@
Method KNN+ FPR95 $\downarrow$ AUROC $\uparrow$ HiMS-Min 37.50 91.00
@@ 8.4 System Accuracy
  • Building a database using all train set images with seen labels, and doing inference only on a test set of images with seen labels where our system scores a 24.42% in total accuracy.
  • Building a database using all train set seen images and newly labeled (previously unseen) examples of images flagged as OOD by the OOD detector module. Finally, we do inference only on images with unseen labels. Our experiments resulted in a 12.50% total accuracy.
In addition to the total accuracy of Veri-Car, we report accuracy when allowing for a 0.2 CER rate, and the accuracy when omitting the License Plate Recognition model results all together. As discussed previously, we expect the license plate model performance to change when finetuning on the license plate layout of a particular country, which would eventually adjust the presented performance metrics. As such, we provide a lower limit (Total Accuracy) and upper limit (Total Accuracy no LPs) to the final expected performance.
Results in Table 18 show an improvement of 21.42% and 5.68% in accuracy when CER $<$0.2 in Seen / Seen and Seen & Unseen / Unseen experiments respectively. At the same time, color prediction showed high synergy with make, type, model, and year. This is due to the shared representation learned by backbone, which allows for correct and incorrect predictions to be aligned with each other.
@@
Database Set / Query Set Total Accuracy Total Accuracy Total Accuracy with CER $<$0.2 no LPs Seen / Seen 24.42 45.84 83.72 Seen & Unseen / Unseen 12.50 18.18 81.25
To the extent of our knowledge, there does not exist an open source system which includes make, model, type, year, color, license plate, and out of distribution detection capabilities, so we cannot benchmark our full system against anything else. There are proprietary models which perform these tasks, but we cannot test against such methods either. That being said, we are still able to identify some real advantages and disadvantages of Veri-Car. Veri-Car is able to reject input samples from classes it does not have in its database, increasing the trustworthiness of its predictions. At the same time, Veri-Car allows for extension of the database beyond its original training set classes, such an extension allows for classification of new classes without retraining. However, since Veri-Car is a retrieval-based system, and thus inference will take longer than standard classifiers with similar number of parameters.
8.5 System Runtime
Finally, we will explore the runtime of the system and the comparison between the version of system using the combined Make, Type, Model, Year and Color model and using a standalone models for both Make, Type, Model and Year and Color. Table 19 shows that combining the standalone models into a single model yields a speed up of 1.82x and of 1.48x when testing the Make, Type, Model, Year and Color (MTMYC) extraction only and the full system, respectively. And only at a small cost of embedding quality and classification power as seen in Section 5.8. Inference times presented are the average over 10 runs on a machine with a single A10 GPU and 16 CPU cores.
Table 19: System runtime comparison Combined Model vs Standalone Model.
System Version MTMYC Only Full System Veri-Car-Combined 22.25 $\pm$ 0.26 imgs/s 13.21 $\pm$ 0.19 imgs/s Veri-Car-Standalone 12.21 $\pm$ 0.02 imgs/s 8.90 $\pm$ 0.01 imgs/s
@@ 9 Conclusion
This paper introduces Veri-Car, a state-of-the-art open-world vehicle information retrieval system system. Veri-Car is able to accurately identify make, type, model, year, and color not only for models it was trained on, but also for up-to 36 additional models without retraining. To address the lack of color labels in the Stanford dataset, which already includes make, type, model, and year labels, we have used an auto-labeling procedure to add color labels based on the annotated VCoR dataset. Experiments show that a retrieval approach to color recognition achieves better results than the standard classification approach. Veri-Car also includes a license plate retrieval component, which leverages synthetic license plates to accurately detect, and recognize both single country and multinational plates with higher accuracy than the baseline model. Moreover, OOD detection results show that HiMS-Min, our new hierarchical version of MS loss, demonstrates not only superior detection capabilities but also FPR@95 robustness when new classes are added into the database. At the same time, experiments show that acquiring 12 samples per-new class is enough to not only accurately infer its characteristics, but to classify it as ID as well.
There are some limitations of this work, which we acknowledge. Our study is limited to the use of open source datasets. Furthermore, like other open-world problems, Veri-Car would require retraining if a substantial number of new samples are introduced.
Future work involves the application of this framework to other industries such as retail, e-commerce, and healthcare. Should a more comprehensive dataset become available, future work would include updating the current models to include a wider variety of data. In addition, future work includes exploring an end-to-end solution which enables license plate, make, type, model, year, and color extraction in a single shot, thereby streamlining the data processing pipeline and enhancing the efficiency of vehicle identification systems.
@@
© 2025 JPMorgan Chase & Co. All rights reserved
Appendix A Model Prediction Visualizations
This section presents visual representations of Make, Type Model, Year & Color Retrieval model including images with correctly and incorrectly predicted labels.
Figure 12: Images where all labels are correctly predicted by Make, Type, Model, Year & Color Retrieval model.
Figure 13: Images where at least one label is incorrectly predicted by Make, Type, Model, Year & Color Retrieval model. GT stands for ground truth and P for predicted label.
Appendix B Bad Quality Examples - Stanford Cars Dataset
This section presents visual representations of low-quality images from the Stanford Car dataset. Many of these low-quality images are affected by adverse weather conditions, such as snow and rain.
Figure 14: Bad quality and bad weather images.
Appendix C Veri-Car: Two Model System Overview
In the main part of the paper, we have introduced a model which can simultaneously detect a car’s make, type, model, year, and color. As discussed, we do not have a single dataset which contains all of these labels, so we used an auto-labeling procedure to add color labels to our main dataset. An alternative to this approach is to separately detect the make, type, model, and year and the color components and to use only manually labeled data for each model. In the main part of the paper, we have chosen to present the unified model as it is simpler and also faster at inference time. For completeness, we outline two model Veri-Car variant here. Many model components remain the same as the single model system so, where relevant, we refer the reader back to the main part of the paper.
Figure 15 shows a high-level view of the two model Veri-Car overview. In this case, Veri-Car is composed of 3 modules: License Plate Retrieval, Make-Model-Type-Year Retrieval, and Color Retrieval. During inference, a new image of a car is provided as input. As in the single model Veri-Car, we first detect the car and crop it. However, we do OOD detection using the make-model-type-year model only. If the car in the image is considered ID, then the Color and LPR modules are used. The next sections describe in detail the performance of each of the Make-Model-Type-Year and Color Retrieval modules independently.
Figure 15: Veri-Car Architecture.
C.1 Make, Type, Model, Year Retrieval model
C.1.1 Dataset
The primary dataset used for the Make, Type, Model, Year Retrieval component is the Stanford Cars 196 Dataset [2], as described in Section 5.1.
C.1.2 Architecture
We use the same architecture as in the Make, Type, Model, Year, and Color Retrieval model, as described in Section 5.2, but we use only one linear layer on top of the image encoder.
C.1.3 Model Configurations
We use the same model configurations as the single model, as described in Section 5.5, but omit cross-entropy.
C.1.4 Training details
The training details are the same as for the full model, as described in Section 5.6.
C.2 Color Retrieval
In this subsection, we give the details on the training configuration, architecture and dataset used to train the Color Retrieval model.
C.2.1 Dataset
The color retrieval task was trained independently of the rest of the system since we found no open source datasets that included all relevant car characteristics. For this reason, the color label cannot be integrated into the hierarchy. Instead, the Color Retrieval model is trained using the Kaggle Vehicle Color Recognition (VCoR) dataset [1], which comprises 15 color classes and 10.5K images.
C.2.2 Architecture
We use the same architecture as the Make, Type, Model, Year retrieval model.
C.2.3 Model Configurations
In order to properly asses our model and choice of loss function we benchmark MS loss against Cross Entropy and Rachmadi et al. [27] Two-Stream Cross Entropy approach:
  • MS: Original MS loss (Equation 1) with a linear layer of size 128 placed on top of the backbone.
C.2.4 Training Details
We process the images the same way as in the Make, Type, Model, Year Retrieval model, but crop a slightly larger area, at least 66%, to prevent cropping out the body of the car. Our training procedure is the same, though this module includes the addition of Cross Entropy which cycles through a minimum of $5\times 10^{-6}$ and maximum of $1\times 10^{-4}$ for 50 epochs.
C.3 Evaluation Metrics
We use the same evaluation metrics as the full model, described in Section 5.7.
C.4 Results
C.4.1 Make, Type, Model, Year Retrieval Model
Table 20(a) shows the retrieval performance for Make, Type, Model, Year Model when only seen classes exist in the database. Results show that in this case HiMS-Min outperforms baselines on Precision@1 while coming in second, behind MS, on mAP@R. Table 20(b) shows similar results on a combined dataset that takes into account both seen and unseen classes. Results in Table 20(b) show that HiMS-Min and MS can produce better separation of seen and unseen classes than other methods. Tables 20(c) and 20(d) show the breakdown of the combined performance found in Table 20(b). Here we can see that HiMS-Min beats most methods in both metrics in the unseen set, except HiMS-Max which ties it in Precision@1, while beating baselines on Precision@1 in the seen set. This points towards HiMS-Min being able to find general features that distinguish between car models even if they were not seen during training. We believe these effects are caused by creating relationships at all levels of the hierarchy, but not inflating the importance at any one level. Furthermore, Figure 16(a) shows the progressive fall in the performance of the model as we add unseen classes to the database and test on them. The performance trend of HiMS-Min and MS point towards HiMS-Min being more robust to classes being added into the database.
Method mAP@R Precision@1 HiSupCon 68.24 93.20 MS 94.78 96.13 HiMS-Max 91.65 96.03 HiMS-Min 94.11 96.18
(a) Seen
Method mAP@R Precision@1 HiSupCon 64.30 91.83 MS 86.38 94.46 HiMS-Max 83.60 94.52 HiMS-Min 86.02 94.78
(b) Combined
Method mAP@R Precision@1 HiSupCon 65.18 91.77 MS 92.03 95.11 HiMS-Max 88.73 94.83 HiMS-Min 91.16 95.15
(c) Seen with combined database
Method mAP@R Precision@1 HiSupCon 60.23 92.11 MS 60.23 91.41 HiMS-Max 59.92 93.08 HiMS-Min 62.30 93.08
(d) Unseen with combined database
To properly asses the quality of the model, we study the minimum amount of samples per-class needed in the database to correctly classify an image from an unseen class. Therefore, we subsample our unseen classes, add them into the seen database, and then test retrieval of our unseen query set. Figure 16(b) shows the average result over 100 runs. We can observe that using hierarchical supervision helps MS trained models to be less sensitive to having less samples in the database, with HiMS-Min scoring $82.99\%$ with 8 samples in the database. We believe this is due to the hard positive and negative mining employed by MS loss. Learning only through these samples at all levels reduces the bias introduced by easy, or redundant samples, helping the model focus on the distinctive features at all levels of hierarchy.
As explained in Section 5.4, the model should be able to correctly assign labels at higher levels of the hierarchy to the samples it incorrectly predicted at the last level of the hierarchy. Table 21 shows that as we go up in the hierarchy, models get better at classifying samples that were incorrectly classified at the Year level (lowest level) of the hierarchy. Additionally, we can see that HiMS-Max and HiSupCon outperform HiMS-Min. This was expected due to the lesser importance placed on creating relationships at higher hierarchical levels. Nonetheless, HiMS-Min still makes a small gain, at the Make level of the hierarchy, over MS loss. Along with that, all models performed poorly at the Model level due to the dataset not having a diversity of years associated to each model.
Method Precision@1 Make-Type-Model Make-Type Make HiSupCon 1.10 48.02 78.63 MS 1.16 48.84 77.13 HiMS-Max 1.51 52.08 78.87 HiMS-Min 1.18 46.67 77.25
C.4.2 Color Retrieval Model
Table 22: Color retrieval model results.
Method Color mAP@R Precision@1 Top-1 Acc. MS 89.68 90.55 Cross Entropy 85.31 88.09 90.29 Two-Stream 84.29 87.53 89.52
Figure 17: MS loss color confusion matrix. Each cell reports the percentage of observations falling in the combination of predicted vs true labels.
Appendix D OOD Detection Results
The OOD detection module remains the same structurally as in the single model system, though in this case we only apply the OOD on the Make, Type, Model, and Year component. Therefore, we report the results for the two model Veri-Car, but do not discuss the method since it remains the same as described in Section 6.
Table 23 shows OOD results for models trained with hierarchical losses and MS loss. For KNN+ experiments, we ran an exhaustive search for the best $k$ and landed on $k=1$. We not only want to optimize the FPR, but we would also like to find the minimum amount of samples needed to be added to the ID database in order to recognize new class samples as ID. The search took into account both FPR@95 and the average recall when a single sample of a new class is added to the ID database. While $k=1$ met both objectives for all methods, $k=4$ achieves the lowest FPR@95 for HiMS-Min, and $k=1$ provided the best trade-off between average recall and FPR@95 (Figure 18). Sun et al. [42] showed that datasets with a small number of samples per-class do not need to retrieve many neighbours to detect OOD samples. In our case, the Stanford Cars dataset averages 41 samples per class in the database, which is relatively small, and thus it makes sense that our chosen $k$ is lower than in other datasets. All methods perform better when using KNN+, which could be due to the lack of a sufficient number samples in the ID set to properly create a Gaussian ellipsoid, as required by the Mahalanobis distance [50]. Moreover, Figure 18 shows that HiMS-Min outperforms all baselines in FPR@95 if we were to choose $k=4$. Experiments in Figure 19 not only shows a similar drop-off in performance when a single sample is added into the ID database, but a general decrease in performance as well for all methods. We believe this is due to, as the low mAP@R but high Precision@1 in Table 20(d) suggest, the quality of the clusters is not high leading to samples of unseen being more spread-out throughout the embedding space. Results shown are the average over 10 random runs for all 36 unseen classes. Furthermore, Figure 20(b) shows a simplified version of Figure 19. We can observe that HiSupCon only needs 4 samples to classify $>80\%$ of the samples of the new class as ID, while HiMS-Min and MS need 8 samples to achieve similar performance.
Table 23: OOD results.
Method KNN+ Mahalanobis FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ HiSupCon 63.14 83.38 91.38 61.26 MS 28.44 92.98 83.89 72.32 HiMS-Max 35.18 91.23 81.79 73.97 HiMS-Min 28.72 93.10 85.47 70.80

### v4

@@ 6.2 License Plate Detection
Veri-Car uses an open-source YOLOv5 model fine-tuned on the LPD task to detect and crop the license plates from car images Face [5]. The model can be further fine-tuned for a specific dataset. However, in the interest of maintaining generalizability and due to the high accuracy of the original model, we decided to use the model out of the box.
@@ 6.3 License Plate Recognition
For LPR, Veri-Car uses an open-source TrOCR model. This model was first fine-tuned on the IAM handwriting dataset as it provided a good warm start relative to the pre-trained only model which was trained on a combination of textlines from PDF documents and synthetically generated handwritten text images [6, 7, 8]. Additional fine-tuning was performed on 35771 generated synthetic license plate images. We will discuss the process for generating synthetic license plate images in Section 6.5.2.
@@ 6.6.2 License Plate Recognition
The results in terms of percentage correct and CER are shown in Tables 7 and 8 respectively. As expected, the fine-tuned TrOCR model performs well on the synthetic license plates and less well on the CCPD license plates, while Tesseract fails on both. It is worth noting that in some subsets of the CCPD data, the model trained only on synthetic data performs decently, demonstrating the power of synthetic data. The TrOCR with CCPD model and the TrOCR with CCPD + Synthetic model perform very well on CCPD license plates, with the latter slightly edging out the former in most cases likely due to the larger amount of data. Given fine-tuning on just 1% of the data, however, TrOCR with CCPD is able to achieve nearly comparable performance, demonstrating that pre-training with synthetic data is a viable option in cases where there are few country specific labels. The TrOCR with CCPD + Synthetic mdel also performs well on just the synthetic data while the performance of TrOCR with CCPD in this case begins to drop off, indicating that during the fine-tuning process the model begins to learn more country specific information. Finally, the results also show that an OCR without LPR specific training, such as Tesseract, fails in the LPR setting.
@@ 6.6.3 End-to-end Evaluation
The end-to-end results shown in Tables 9 and 10 depict RPnet’s failure to capture license plates from countries on which it was not trained, even when the data has been filtered to include only license plates which fit the pattern of the training data. On the other hand, Veri-Car’s YOLO+TrOCR model, which is fine-tuned on synthetic data only, offers stronger generalizability, outperforming RPnet by a wide margin on the Stanford Cars dataset which includes license plates from a range of countries. The results further underscore the effectiveness of Veri-Car’s two additional TrOCR variants, both of which often surpass PRNet on the CCPD data. This shows that synthetic data can be valuable not only in a multinational context, but also when applied during pre-training for a country-specific detector. Similar to the LPR results, TrOCR with CCPD + Synthetic slightly outperforms TrOCR with CCPD, likely due to the increased volume of training data. Notably, within the green subset of CCPD plates, which represent images of green license plates while the training data contains only blue plates, TrOCR with CCPD significantly outperforms all other models. This suggests that the use of synthetic data is especially effective in accurately recognizing license plate numbers, despite the color variation in these plates.
@@ 7 Full System Evaluation
In this section we evaluate the effectiveness of the system using a subset of 101 samples of the Stanford Cars dataset [51]. We use the metrics already presented in each component of Veri-Car. As discussed in previous sections we struggled to find a training dataset with all required labels. To overcome this issue we use a dataset with complete labels for everything, but the car color. We asked three researchers to label the images, and used a voting system to decide the final labels. Our results are summarized in Table 11. This new dataset includes 85 images of cars with labels seen during training and 16 images with unseen labels, never seen during training. The results in Table 11 summarize the performance of the each component of Veri-Car for all images corresponding to seen label instances. In this section we also present the results of the Out-of-Distribution detection model on the unseen set of images and summarize the overall performance of the system in correctly classifying all labels of an image on both the seen and unseen sets independently.
@@
Model Metric Result Color Retrieval MAP@R 75.08 Prec@1 78.14 Color Retrieval (Popular Colors) MAP@R 85.17 Prec@1 86.39 Make, Type, Model, Year retrieval MAP@R 83.43 Prec@1 88.21 License Plate Recognition Percent Correct 28.43 CER 21.50 License Plate Recognition (CER <0.2) Percent Correct 49.15 CER 7.71
7.1 Color Retrieval Model
The Color Retrieval model achieved a MAP@R of 75.08 and a Prec@1 of 78.14. Through the labelling process, we observed that labeling certain colors can be ambiguous, particularly for less common colors like beige, tan and purple, potentially resulting in labeling errors. Recognizing the potential influence of less popular colors on model performance, we conducted an additional experiment focusing on only the popular color labels, including black, blue, grey, red, silver, and white. The results of this experiment showed a notable performance improvement of approximately 10% in MAP@R and 8% in Prec@1. Given the ambiguity in labeling these less common colors, along with the observed performance improvement on the subset of most popular colors, we conclude that the model’s performance could be enhanced by training with additional images of underrepresented colors, if such data becomes available. Increasing the number of images for these car colors would likely reduce the impact of labeling errors, allowing their distribution to reach a steady state. This, in turn, could help the model learn a more accurate representation of these colors and their relationship to similar hues in the dataset. At the same time, colors under different lighting conditions can look like other colors (e.g. brow, beige, tan and gold), so additional labels which include lighting conditions could further improve color recognition models.
7.2 Make, Type, Model, Year Retrieval Model
The model responsible for predicting car characteristics, including manufacturer, model, type, and class, was evaluated on the same test set. The model demonstrated robust performance, achieving a MAP@R of 83.43 and a Prec@1 of 88.21.
@@ 7.4 OOD Detection
The OOD detection model results are reported in Table 12. Out of the 102 images in the test set, 16 were not seen by the model during training. The OOD detection model has an FPR95 OF 6.25% and AUROC of 94.84%, highlighting the success of the OOD detection model in flagging unseen samples.
@@
Method KNN+ FPR95 $\downarrow$ AUROC $\uparrow$ HiMS-Min 6.25 94.84
@@ 7.5 System Accuracy
  • Building a database using all train set images with seen labels, and doing inference only on a test set of images with seen labels
  • Building a database using all train set seen images and newly labeled (previously unseen) examples of images flagged as OOD by the OOD detector model. Finally, we inference only on images with unseen labels.
In addition to the total accuracy of the system, we report accuracy when allowing for a 0.2 CER rate, and the accuracy when omitting the License Plate Recognition model results all together. As discussed previously, we expect the license plate model performance to change when fine tuned on the license plate layout of a particular country, which would eventually adjust the presented performance metrics. As such, we provide a lower limit (Total Accuracy) and upper limit (Total Accuracy no LPs) to the final expected performance.
@@
Database Set / Query Set Total Accuracy Total Accuracy Total Accuracy with CER $<$0.2 no LPs Seen / Seen 30.23 52.33 90.70 Seen & Unseen / Unseen 6.25 56.35 87.50
@@ 8 Conclusion
This paper introduces Veri-Car, a state-of-the-art open-world vehicle information retrieval system system. Veri-Car is able to accurately identify make, type, model, year and color of not only models it was trained on, but up-to 36 additional models without retraining. Experiments show that a retrieval approach to color recognition achieves better results than the standard classification approach. Veri-Car also includes a license plate retrieval component, which leverages synthetic license plates to accurately detect, and recognize both single country and multinational plates with higher accuracy than the baseline model. Moreover, OOD detection results show FPR@95 robustness when new classes are added into the database. At the same time, experiments show that acquiring 8 samples per-new class is enough to not only accurately infer its characteristics, but to classify it as ID as well.
@@
© 2024 JPMorgan Chase & Co. All rights reserved